OpenStudy (anonymous):

What is the height of a triangle that has an area of 60 yd2 and a base with a length of 12 yd? A. 0.1 yd B. 2.5 yd C. 5 yd D. 10 yd

OpenStudy (anonymous):

Does the triangle have a specified type such as right angle or equilateral?

OpenStudy (anonymous):

that's all the question said, no picture as well.

OpenStudy (anonymous):

Ok let me think then

OpenStudy (anonymous):

Area = 1/2 x base x perpendicular ht

OpenStudy (anonymous):

THE IS SUPPOSED TO BE A EXPONENT RIGHT? SORRY CAPS LOCK

OpenStudy (radar):

Thats all you need, use Area A=1/2 b h

OpenStudy (radar):

Area is in "square" units.

OpenStudy (anonymous):

ohhh ok!!!

OpenStudy (anonymous):

let me check then and see

OpenStudy (anonymous):

so 1/2*B*H?

OpenStudy (radar):

Yes 60 sq yd = 1/2 12 yd h(yd) 60 yd^2 = 6 h (yd) h = ?? yd.

OpenStudy (anonymous):

10 is my anwser

OpenStudy (radar):

I see 10 as a choice. Make the right choice. Good luck.

OpenStudy (anonymous):

10*12*1/2=60

OpenStudy (radar):

Good proof.

OpenStudy (anonymous):

did i do it right?

OpenStudy (radar):

Your procedure is fine, and the answer unit is yards.

OpenStudy (anonymous):

do i have to square root both sides?

OpenStudy (radar):

No, your work on this problem is fiished.

OpenStudy (anonymous):

ohh so i did it right?

OpenStudy (anonymous):

thanks Radar!~

OpenStudy (radar):

Yes, excellent.

OpenStudy (radar):

You're welcome. Good luck in your studies.